White gold" alloy contains 90% gold and 10% silver. \Red gold" alloy contains 80% gold and 20% copper. How much \white gold" and how much \red gold" should we add to 60g of the alloy containing 45% gold, 50% silver and 5% copper, in order to obtain alloy with 70% gold, 15% silver and 15% copper concentration?
1
Expert's answer
2020-10-12T15:14:32-0400

Let White gold = W

Let Red gold = R

Let gold = g

Let silver = s

Let copper = c


"\therefore" W = 0.9g + 0.1s

"\therefore" R = 0.8g + 0.2c


"xg(W) + yg(R) + 60g(0.45g+0.5s+0.05c) = (x+y+60)g\times(0.7g+0.15s+0.15c)"

where x = mass of White gold; and

where y = mass of Red gold


"x(0.9g+0.1s) + y(0.8g+0.2c) +60(0.45g+0.5s+0.05c) = (x+y+60)(0.7g+0.15s+0.15c)"


"0.9gx + 0.1sx + 0.8gy + 0.2cy + 27g + 30s + 3c = 0.7gx + 0.7gy + 42g + 0.15sx + 0.15sy + 9s + 0.15cx + 0.15cy + 9c"


collecting like terms;

"0.9gx + 0.8gy + 27g + 0.1sx + 30s + 0.2cy + 3c = 0.7gx + 0.7gy + 42g + 0.15sx + 0.15sy + 9s + 0.15cx + 0.15cy + 9c"


"g(0.9x + 0.8y + 27) + s(0.1x + 30) + c(0.2y + 3) = g(0.7x + 0.7y +42) +s(0.15x + 0.15y + 9) + c(0.15x +0.15y +9)"


Comparing both sides

Firstly;

"g(0.9x + 0.8y + 27) = g( 0.7x + 0.7y + 42)"

"0.9x + 0.8y + 27 = 0.7x + 0.7y + 42"

"0.2x + 0.1y = 15" "(\times10)"

"2x + y = 150" ---------(i)


Secondly;

"s(0.1x + 30) = s(0.15x + 0.15y +9)"

"0.1x + 30 = 0.15x + 0.15y +9"

"-0.05x - 0.15y = -21" "(\times-20)"

"x + 3y = 420" ---------(ii)


Thirdly;

"c(0.2y + 3) = c(0.15x + 0.15y +9)"

"0.2y + 3 = 0.15x + 0.15y +9"

"0.05y - 0.15x = 6" "(\times-20)"

"y - 3x = 120" ---------(iii)


Making equation (i) and equation (iii) into simultaneous equations

"2x + y = 150" ---------(i)

"y - 3x = 120" ---------(iii)

Subtracting equation (ii) from equation (i)

"2x-(-3x) +y -y = 150 -120"

"5x = 30"

"\therefore x = 6"


Substituting the value of x into equation (ii)

"y - 3x = 120"

"y = 120 + 3x"

"y = 120 + 3(6)"

"y = 138"


"\therefore" 6g of White gold and 138g of Red gold should be added